pooja195 (pooja195):

.Rationalize the denominators of the following expressions and simplify if possible.

OpenStudy (i<3u):

Spam?

OpenStudy (alexandervonhumboldt2):

no

pooja195 (pooja195):

Totally. \[\huge~\rm~\frac{ 4 }{ \sqrt{3}-\sqrt{2} }\]

pooja195 (pooja195):

Rationalize the denominators of the following expressions and simplify if possible.

OpenStudy (alexandervonhumboldt2):

ok this is easy

OpenStudy (alexandervonhumboldt2):

\[\frac{ 4 }{ \sqrt3 - \sqrt2 }=\frac{ 4 }{ \sqrt3 - \sqrt2 }*\frac{ \sqrt3 + \sqrt2 }{ \sqrt3 +\sqrt2 }\]

OpenStudy (alexandervonhumboldt2):

\[\frac{ 4*(\sqrt3 + \sqrt2) }{ 3-2 }\]

OpenStudy (alexandervonhumboldt2):

now you can simpify :) (a-b)*(a+b)=a^2-b^2

OpenStudy (astrophysics):

conjugateeeee nation

OpenStudy (astrophysics):

https://www.mathsisfun.com/algebra/conjugate.html

pooja195 (pooja195):

Isnt the bottom supposed to me \[\huge~\rm~\sqrt{3^2}-\sqrt{2^2}\] Thanks astro for the link

pooja195 (pooja195):

*be not me

OpenStudy (alexandervonhumboldt2):

i just used the difference of squares @pooja195

OpenStudy (astrophysics):

\[\left( \sqrt{3} \right)^2-\left( \sqrt{2} \right)^2 = 3-2\]

OpenStudy (alexandervonhumboldt2):

you just written it in another form

OpenStudy (alexandervonhumboldt2):

3-2=(sqrt3-sqrt2)*(sqrt3+sqrt2)

OpenStudy (astrophysics):

Remember \[\sqrt{x} = x^{1/2}\] so when you sqaure it you get \[x^{2/2} = x\]

OpenStudy (astrophysics):

Difference of squares works to :)

pooja195 (pooja195):

Would my final answer be \[\huge~\rm~4(\sqrt{3}+\sqrt{2})\]

OpenStudy (alexandervonhumboldt2):

yes, but maybe you need to distribute it?

OpenStudy (astrophysics):

\[\huge \checkmark\]

pooja195 (pooja195):

I need to distribute???

OpenStudy (astrophysics):

Nope, you're good

OpenStudy (alexandervonhumboldt2):

no just asked

pooja195 (pooja195):

lol ok thanks guys ^-^

OpenStudy (astrophysics):

The point of these problems is just to get you used to using the conjugate

pooja195 (pooja195):

lol I suck at these i find them difficult because im not used to using the conjugate :P

OpenStudy (astrophysics):

Just read that link I sent you it's barely a page long should clear things up, it's just a change of sign

pooja195 (pooja195):

I see thanks for the link its actually very helpful =)

OpenStudy (alexandervonhumboldt2):

yeah, because if you use conjugate it is simpler because of difference of squares. np :D :) :\ :O

pooja195 (pooja195):

lol I see yeah thanks :-)